OpenStudy (anonymous):

which fraction is NOT equivalent to 4\12? A. 1\3 B. 12\30 C. 8\24 D. 16\48

OpenStudy (calculusxy):

lets try the fractions out. okay so 4/4=1 12/4=3 = 1/3 (so not A) 12/12= 1 but 30/12=2.5 (B is on hand) 8/4=2 so 4*8=2 and 12*2=24 = 8/24 16/4=4 so 4*4=16 and 12*4=48 = 16*48